# Source Code Vulnerability Detection: Combining Code Language Models and Code Property Graphs

### v1

@@ IV-E Joint Training of CodeBERT and GGNN
In our joint training approach, we optimize the parameters of both CodeBERT and GGNN, leveraging the complementary strengths of each model—CodeBERT’s contextual understanding and GGNN’s relational insights—to improve the model’s performance in detecting code vulnerabilities. This joint optimization strategy is implemented through the use of cross-entropy loss across code graph nodes, allowing for the simultaneous optimization of parameters for CodeBERT and GGNN. The formulated loss function can be depicted as:
@@
Figure 4: Detection accuracy for the top-30 high-frequency CWE vulnerability types in DiverseVul.
In this training process, CodeBERT updates the node embeddings with each iteration, thereby gradually improving the complementary advantages of both CodeBERT and GGNN.
IV-F Interpolating Predictions
In the previous step, we implicitly combine CodeBERT and GGNN by utilizing CodeBERT to generate the node embeddings for GGNN. Here, we further explicitly combine the benefits of pre-training and graph-based approaches by leveraging interpolation predictions. Specifically, we introduce an auxiliary classifier that operates directly on CodeBERT embeddings by feeding code embeddings $E$ into a dense layer with softmax activation. Ultimately, we perform a linear interpolation [31] of the predictions from Vul-LMGNN and CodeBERT, which is expressed as follows:
@@ IV-F Interpolating Predictions
The parameter $\lambda$ controls the trade-off between the two objectives. A value of $\lambda=1$ signifies the exclusive use of the full Vul-LMGNN model, whereas $\lambda=0$ indicates reliance solely on the CodeBERT module. When $\lambda$ is within the range $(0,1)$, it allows for a balanced integration of the predictions from both models. The fine-tuned CodeBERT model regulated and optimized the input graph for the GGNN. Subsequently, an interpolation prediction facilitated an appropriate trade-off between the graph model and sequence model detection results, yielding outstanding detection outcomes.
V Experiments results
@@
    RQ1: How does our Vul-LMGNN performance compare with other learning-based methods for vulnerability identification?
@@
TABLE II: Performance metrics of various models on datasets with specific CWEs.
Dataset Model ACC (%) P (%) R (%) F1 (%)
DiverseVul Bert 91.99 27.95 13.09 17.83
CodeBert 92.40 28.26 20.02 23.44
GraphCodeBert 92.96 31.14 16.30 21.40
TextCNN 92.16 10.25 9.82 10.03
TextGCN 91.50 15.66 11.50 13.27
Devign(AST) 70.21 9.35 9.22 9.28
Our 93.06 32.21 18.54 23.54
Draper VDSIC Bert 79.41 81.86 75.97 78.80
CodeBert 83.13 86.13 78.97 82.39
GraphCodeBert 83.98 84.74 83.17 83.95
TextCNN 66.54 65.36 70.55 67.86
TextGCN 67.55 67.66 67.63 67.64
Devign(AST) 59.30 58.84 68.93 63.49
Our 84.38 87.37 80.64 83.87
The experiments were executed on single NVIDIA A100 80GB GPU. The system specifications comprised NVIDIA driver version 525.85.12 and CUDA version 11.8. The software environment was configured with Python 3.10.13 and torch 2.2.0.
In our comparative analysis, Vul-LMGNN is benchmarked against the latest state-of-the-art detection models. This includes Transformer-based models: Bert [32], CodeBert, and GraphCodeBert; GNN-based models: TextGCN [33] and Devign; As well as the CNN-based model TextCNN [34]. For computational efficiency, functions with a node size exceeding 500 in the CPG were excluded from our analysis. In terms of our model’s configuration, the learning rate and batch size were set to $41e-$ 4 and 64, respectively. The training was conducted over 20 epochs with an early stopping criterion triggered if no further optimization in performance. Specifically for the Devign model, AST was employed for the code graph representation. Given the absence of disclosed hyperparameters, we endeavored to replicate their methodology to the best of our ability. The following are the details of the baseline:
@@
    Bert: A powerful pre-trained language model developed by Google, widely used for natural language understanding tasks.
  • $\bullet$
    CodeBERT: A language model specifically fine-tuned for code-related tasks, including code summarization and code completion.
  • $\bullet$
    GraphCodeBERT: A pre-trained programming language model, expanding upon CodeBERT to integrate code data flow information into the training objective.
  • $\bullet$
    TextCNN: A CNN architecture from the field of natural language processing, widely used in code vulnerability detection [35, 36].
  • $\bullet$
    TextGCN: TextGCN: An advanced method for learning the graph representations from text, showcasing exceptional performance in code-related tasks.
  • $\bullet$
    Devign: A gated GNN-based model, which takes a code property graph as input and employs 1-D convolutional pooling to make predictions.
V-A Comparison with Baselines (RQ1)
To evaluate the performance of Vul-LMGNN on code vulnerability detection, we executed an extensive comparative analysis against six baseline models utilizing the four datasets delineated in Table 1. The experimental results are systematically presented in Table II and III.
We initially tested the Vul-LMGNN on datasets categorized by specific CWEs and analyzed its capability to recognize these CWEs within the test set. In terms of accuracy, precision, and F1 score, Vul-LMGNN outperformed all baseline models. Specifically, within the DiverseVul dataset, our model achieved an accuracy of 93.06% and an F1-score of 23.54%. In the balanced version of the VDSIC dataset, an accuracy of 84.38% was attained.
As shown in Figure 4, among the top 30 most frequently occurring CWEs in the test set, our model achieved a highest accuracy rate of 50% CWEs. It can be observed that for some CWEs, the recognition accuracy of the model is generally low, such as CWE-310 (Cryptographic Issues) and CWE-189 (Numeric Errors), while for another subset of CWEs, there are high recognition accuracy rates, such as CWE-134 (Controlled Format String) and CWE-770 (Allocation of Resources Without Limits or Throttling).
Among the baseline sequence-based detection models, CodeBert and GraphCodeBert showcased superior detection capabilities owing to their programming language pre-training tasks, despite not containing C/C++ programs in their pre-training datasets. As a component of our model, CodeBert attained accuracies of 92.40% and 83.13%, and precisions of 28.26% and 86.13%, respectively. These values were marginally lower by 0.66% and 1.25%, and 3.95% and 1.67%, compared to our model. GraphCodeBert, with further integration of data flow information, outperformed CodeBert, reducing the precision gap with our model to 1.07%, although the recall gap widened to 2.24%.
@@ V-A Comparison with Baselines (RQ1)
TABLE III: Performance metrics of various models on datasets with no specific CWEs.
Dataset Model ACC (%) P (%) R (%) F1 (%)
Devign Bert 60.58 57.67 54.64 56.11
CodeBert 63.93 60.30 63.00 61.62
GraphCodeBert 64.80 64.37 54.38 58.96
TextCNN 60.38 59.03 57.72 58.37
TextGCN 60.47 60.87 58.58 59.70
Devign(AST) 57.66 56.96 56.25 56.60
Our 65.70 64.53 56.34 60.16
ReVeal Bert 86.88 32.70 40.13 36.04
CodeBert 88.64 38.26 38.13 38.19
GraphCodeBert 89.25 41.67 41.81 41.74
TextCNN 85.43 26.32 20.33 22.94
TextGCN 87.25 24.61 17.85 20.69
Devign(AST) 65.47 17.38 18.09 17.72
Our 90.80 57.09 46.45 51.22
In the realm of graph-based detection models, TextGCN, while performing well in text classification, showed mediocre results in code vulnerability detection experiments, achieving only a 91.50% accuracy rate on DiverseVul, with precision and recall at 15.66% and 11.50%, respectively. This may be due to TextGCN’s focus on word co-occurrence, lacking structural code information. The AST version of Devign, which incorporates control flow and data flow information and uses Word2Vec along with the average of tokens for node vector representation, performed poorly with an accuracy of only 70.21%, a gap of 14.26% from our F1 score. This could be attributed to the neglect of local semantic information of code within the node. These disparities were more pronounced on the VDSIC dataset, with gaps in accuracy reaching 16.83% and 25.08% compared to our model, respectively.
The performance disparity on the other two datasets without specific CWEs is similar to those observed in the previous CWE-specific evaluations, as illustrated in Table III.
Overall, transformer-based models demonstrated better detection effectiveness, as the embedding layer of transformers can implicitly capture vulnerability-related signals from the source code. In contrast, methods like Word2Vec, trained by predicting adjacent tokens, extract contextual information that may not be effective for vulnerability detection. Vul-LMGNN, utilizing a programming language (PL) model and GNNs, preserves the sequential information in the code and better incorporates its inherent information.
@@
Figure 5: Accuracy of Vul-LMGNN when varying trade-off parameter on partial DiverseVul dataset.
@@
Figure 6: Precision, recall and f1 of Vul-LMGNN when varying trade-off parameter on partial DiverseVul dataset.
V-B Impact of the Tradeoff Parameter (RQ2)

### v2

@@ 3.6.1 Implicit Stage
In our joint training approach, we optimize the parameters of both CodeLM and GGNN, leveraging the complementary strengths of each model—CodeLM’s contextual understanding and GGNN’s relational insights—to improve the model’s performance in detecting code vulnerabilities. This joint optimization strategy is implemented through the use of cross-entropy loss across code graph nodes, allowing for the simultaneous optimization of parameters for CodeLM and GGNN. The formulated loss function can be depicted as:
@@
In this training process, CodeLM updates the node embeddings with each iteration, thereby gradually improving the complementary advantages of both CodeLM and GGNN.
3.6.2 Explicit Stage
In the previous step, we implicitly combine CodeLM and GGNN by utilizing CodeLM to generate the node embeddings for GGNN. Here, we further explicitly combine the benefits of pre-training and graph-based approaches by leveraging interpolation predictions. Specifically, we introduce an auxiliary classifier that operates directly on CodeLM embeddings by feeding code embeddings $E$ into a dense layer with softmax activation. Ultimately, we perform a linear interpolation [42] of the predictions from Vul-LMGNNs and CodeLM, which is expressed as follows:
@@ 3.6.2 Explicit Stage
The parameter $\lambda$ controls the trade-off between the two objectives. A value of $\lambda=1$ signifies the exclusive use of the full Vul-LMGNNs model, whereas $\lambda=0$ indicates reliance solely on the CodeBERT module. When $\lambda$ is within the range $(0,1)$, it allows for a balanced integration of the predictions from both models. The fine-tuned CodeLM model regulated and optimized the input graph for the GGNN. Subsequently, an interpolation prediction facilitated an appropriate trade-off between the graph model and sequence model detection results, yielding outstanding detection outcomes.
Table 1: Summary of Datasets
Dataset #Vulnerable #Non-Vul Source CWEs
DiverseVul 18,945 330,492 Snyk, Bugzilla 150
Devign 11,888 14,149 Github N/A
VDSIC 82,411 119,1955 GitHub, Debian 4
ReVeal 1664 16,505 Chrome, Debian N/A
4 Dataset Review
To evaluate our proposed code vulnerability detection method and other baseline methods, it is imperative to possess a substantial quantity of both vulnerable and non-vulnerable source code, spanning a diverse range of vulnerabilities. In this paper, we have selected four public code vulnerability datasets, which include three widely-used popular datasets and one newly released comprehensive dataset. We have summarized the distribution of positive and negative samples and sources of the datasets, as well as whether they distinguish specific types of vulnerability, as shown in Table 1.
5 Experiments results
@@
    RQ1: How does our Vul-LMGNNs performance compare with other learning-based methods for vulnerability identification?
@@
    RQ5: Does online knowledge distillation enhance model performance? How does the number of student models alter performance?
The experiments were executed on a single NVIDIA A100 80GB GPU. The system specifications comprised NVIDIA driver version 525.85.12 and CUDA version 11.8. The software environment was configured with Python 3.10.13 and torch 2.2.0.
@@
Table 2: Performance metrics of various models on datasets with specific CWEs.Note: C-B: CodeBERT, GC-B: GraphCodeBERT, C-T5s: CodeT5-Small, C-T5b: CodeT5-Base.
DiverseVul Dataset Model ACC (%) P (%) R (%) F1 (%) Bert 91.99 27.95 13.09 17.83 CodeBert 92.40 28.26 20.02 23.44 GraphCodeBert 92.96 31.14 16.30 21.40 CodeT5-Small 94.27 43.00 17.46 24.83 CodeT5-Base 94.24 42.96 19.05 26.40 TextCNN 92.16 10.25 9.82 10.03 TextGCN 91.50 15.66 11.50 13.27 Devign(AST) 70.21 9.35 9.22 9.28 VulBERTa-MLP 92.14 21.47 16.92 18.92 VulBERTa-CNN 91.33 17.76 16.51 17.11 TokenCNN 90.54 11.21 10.76 10.98 VulDeePecker 91.76 17.31 13.76 15.33 SySeVR 92.86 27.10 18.76 22.17 CFExplainer 88.86 13.63 19.76 16.13 ReGVD 85.86 10.26 20.16 13.73 VulMAE 88.78 11.21 15.46 13.00 ReVeal 85.88 9.35 18.46 12.42 Vul-LMGNNs(C-B) 93.06 32.21 18.54 23.54 Vul-LMGNNs(GC-B) 93.61 33.73 18.51 23.90 Vul-LMGNNs(C-T5s) 95.04 60.51 20.01 30.43 Vul-LMGNNs(C-T5b) 94.84 55.86 23.01 32.59 Draper VDSIC Dataset Model ACC (%) P (%) R (%) F1 (%) Bert 79.41 81.86 75.97 78.80 CodeBert 83.13 86.13 78.97 82.39 GraphCodeBert 83.98 84.74 83.17 83.95 CodeT5-Small 84.47 84.21 84.70 84.51 CodeT5-Base 85.01 85.90 84.64 85.12 TextCNN 66.54 65.36 70.55 67.86 TextGCN 67.55 67.66 67.63 67.64 Devign(AST) 59.30 58.84 68.93 63.49 VulBERTa-MLP 80.41 83.37 75.97 79.50 VulBERTa-CNN 78.98 79.46 78.17 78.81 TokenCNN 65.05 63.99 68.84 66.33 VulDeePecker 78.15 78.50 77.53 78.01 SySeVR 82.01 83.25 80.15 81.67 CFExplainer 80.01 83.37 74.97 78.95 ReGVD 79.78 80.15 79.17 79.66 VulMAE 77.05 76.12 78.84 77.45 ReVeal 78.15 78.50 77.53 78.01 Vul-LMGNNs(C-B) 84.38 87.37 80.64 83.87 Vul-LMGNNs(GC-B) 84.51 86.36 81.97 84.11 Vul-LMGNNs(C-T5s) 85.68 86.69 84.31 85.48 Vul-LMGNNs(C-T5b) 85.91 86.33 85.33 85.83
In our comparative analysis, we evaluated the performance of Vul-LMGNN against 17 state-of-the-art detection models. This includes 5 Transformer-based models: BERT [45], CodeBERT [15], GraphCodeBERT [16], CodeT5-Small [17], and CodeT5-Base [17]. We also included 2 classical text-based techniques commonly used for vulnerability detection: TextGCN [46] and TextCNN [47]. In addition, we evaluated 10 specialized vulnerability detection algorithms: Devign [31], VulBERTa-MLP [48], VulBERTa-CNN [48], TokenCNN [4], VulDeePecker [8], SySeVR [14], CFExplainer [49], ReGVD [24], VulMAE [50], and ReVEAL [32]. For computational efficiency, functions with a node size exceeding 500 in the CPG were excluded from our analysis. In terms of our model’s configuration, the learning rate and batch size were set to $41e-$ 4 and 64, respectively. The training was conducted over 20 epochs with an early stopping criterion triggered if no further optimization in performance. Specifically for the Devign model, AST was employed for the code graph representation.
To ensure fairness, we used official code repositories for baseline models when available and followed original papers closely where not. All models were trained under identical conditions, including hardware, dataset splits, batch sizes, and optimizers. Key parameters were set according to original recommendations or through grid search.
Table 3: Performance metrics of various models on datasets with no specific CWEs.Note: C-B: CodeBERT, GC-B: GraphCodeBERT, C-T5s: CodeT5-Small, C-T5b: CodeT5-Base.
Devign Dataset Model ACC (%) P (%) R (%) F1 (%) Bert 60.58 57.67 54.64 56.11 CodeBert 63.93 61.03 58.10 59.53 GraphCodeBert 64.80 64.37 54.38 58.96 CodeT5-Small 65.62 64.35 55.39 59.53 CodeT5-Base 65.92 64.27 57.12 60.48 TextCNN 60.38 59.03 57.72 58.37 TextGCN 60.47 60.87 58.58 59.70 Devign(AST) 57.66 56.96 56.25 56.60 VulBERTa-MLP 64.75 62.71 56.22 59.29 VulBERTa-CNN 64.42 63.11 53.12 57.69 TokenCNN 47.63 43.98 45.01 43.97 VulDeePecker 60.36 56.94 55.41 56.39 SySeVR 62.18 60.45 54.05 57.30 CFExplainer 57.66 56.96 56.25 56.60 ReGVD 61.89 60.74 48.20 53.75 VulMAE 62.36 60.47 50.71 55.16 ReVeal 62.38 59.80 53.71 56.59 Vul-LMGNNs(C-B) 65.70 64.53 56.34 60.16 Vul-LMGNNs(GC-B) 66.33 64.73 57.77 61.01 Vul-LMGNNs(C-T5s) 66.77 63.45 64.20 63.82 Vul-LMGNNs(C-T5b) 67.27 64.73 62.20 63.44 ReVeal Dataset Model ACC (%) P (%) R (%) F1 (%) Bert 86.88 32.70 40.13 36.04 CodeBert 88.64 38.26 38.13 38.19 GraphCodeBert 89.25 41.67 41.81 41.74 CodeT5-Small 90.50 47.81 40.73 43.99 CodeT5-Base 90.94 50.69 39.76 44.56 TextCNN 85.43 26.32 20.33 22.94 TextGCN 87.25 24.61 17.85 20.69 Devign(AST) 87.49 36.65 31.55 33.91 VulBERTa-MLP 88.48 36.79 35.90 36.34 VulBERTa-CNN 87.64 34.46 38.76 36.48 TokenCNN 73.48 17.47 50.90 26.01 VulDeePecker 89.05 17.68 13.87 15.70 SySeVR 84.22 24.36 40.11 30.25 CFExplainer 85.18 24.21 29.01 26.39 ReGVD 90.63 64.70 14.47 23.65 VulMAE 88.24 30.25 21.76 25.31 ReVeal 85.37 29.90 40.91 33.87 Vul-LMGNNs(C-B) 90.80 57.09 46.45 51.22 Vul-LMGNNs(GC-B) 91.58 55.12 43.41 48.57 Vul-LMGNNs(C-T5s) 90.98 50.76 50.41 50.58 Vul-LMGNNs(C-T5b) 91.68 54.89 51.41 53.09
5.1 Comparison with Baselines (RQ1)
To evaluate the performance of Vul-LMGNNs on code vulnerability detection, we executed an extensive comparative analysis against 17 baseline models utilizing the four datasets delineated in Table 1. The experimental results are systematically presented in Table 2 and 3.
We initially tested the Vul-LMGNNs on datasets categorized by specific CWEs and analyzed their capability to recognize these CWEs within the test set. In terms of accuracy, precision, and F1 score, Vul-LMGNNs outperformed all baseline models. Specifically, any variant of Vul-LMGNNs outperformed all baseline models. Specifically, on the DiverseVul dataset, Vul-LMGNNs with CodeBERT as a component achieved an accuracy of 93.06% and an F1 score of 23.54%; on the balanced version of the VDSIC dataset, they reached an accuracy of 84.38%. Vul-LMGNNs with CodeT5 as a component achieved an accuracy of 95.04% on the DiverseVul dataset, with an F1 score of 32.59%, surpassing the previous best method by 10%.
Vul-LMGNNs outperform other models primarily due to their effective integration of structural information from GNNs and semantic understanding from pre-trained language models. While models like BERT and CodeBERT capture sequential semantics well, they lack the structural insights needed for vulnerability detection in complex codebases. In contrast, graph-based models like TextGCN struggle in our experiments, as they focus on word co-occurrence and miss critical code structure details. By leveraging both semantic and structural features, Vul-LMGNNs achieve superior performance across various metrics, excelling in both precision and recall when detecting vulnerabilities embedded in the code.
Furthermore, extensive experiments demonstrated that code language models trained on massive datasets generally outperform specialized vulnerability detection methods, despite not containing C/C++ programs in their pre-training datasets. Pre-trained LM like BERT, CodeBERT, and CodeT5 variants consistently achieved higher accuracy (91.99% - 94.27%) compared to specialized methods such as VulDeePecker (91.76%), SySeVR (92.86%), and ReVeal (85.88%). CodeT5-Small and CodeT5-Base demonstrated particularly strong performance, benefiting from their larger architectural scale and pre-training, with accuracies of 94.27% and 94.24% respectively, and F1 scores of 24.83% and 26.40%. The other three datasets exhibitd similar trends. These results strongly suggest that the transfer learning capabilities and rich semantic understanding of large-scale pre-trained code language models provide a significant advantage in vulnerability detection tasks. Their ability to capture complex code patterns and contextualized representations appears to be more effective than the feature engineering approaches typically used in specialized vulnerability detection methods.
As shown in Figure 5, we demonstrated the performance of our method on specific vulnerabilities. We compared the competitiveness of our approach with CodeBert and GraphCodeBert. In this comparison, our method is the Vul-LMGNNs with CodeBert as a component, which we refer to as CodeBertGGNN. Among the top 30 most frequently occurring CWEs in the test set, our model achieved the highest accuracy rate on 50% of CWEs. It can be observed that for some CWEs, the recognition accuracy of the model is generally low, such as CWE-310 (Cryptographic Issues) and CWE-189 (Numeric Errors), while for another subset of CWEs, there are high recognition accuracy rates, such as CWE-134 (Controlled Format String) and CWE-770 (Allocation of Resources Without Limits or Throttling).
We analyzed the lower identification rates for CWE-310 and CWE-189 and identified several potential reasons for these challenges:
Regarding CWE-310 (Cryptographic Issues): These vulnerabilities primarily involve encryption problems, which can manifest in cryptographic algorithms, key management, or protocol implementations within the code. Vulnerability detection tools face difficulties in identifying CWE-310 class vulnerabilities for the following reasons: 1). While the model can learn some indicators of cryptographic vulnerabilities from contextual dependencies, certain encryption issues only become apparent under specific runtime conditions. 2). The complexity and subtlety of cryptographic implementations often require domain-specific knowledge that may be beyond the scope of general-purpose vulnerability detection tools.
Concerning CWE-189 (Numeric Errors): The challenges in detecting these vulnerabilities are similar to those of CWE-310: 1). Some numeric errors only manifest under specific input conditions or runtime environments, making it difficult for automated tools to capture these dynamic behaviors. 2). There is a significant human factor involved in CWE-189 vulnerabilities. Developers may introduce counting errors due to mistakes or oversights during coding, which are often subtle and not easily detectable by automated tools.
In both cases, while our model can infer some vulnerability indicators from code context and structure, it may struggle with issues that only become apparent during runtime or require deep domain-specific understanding.
@@ 5.1 Comparison with Baselines (RQ1)
Figure 6: Accuracy of Vul-LMGNNs when varying trade-off parameter on partial DiverseVul dataset.
@@
Figure 7: Precision, recall and f1 of Vul-LMGNNs when varying trade-off parameter on partial DiverseVul dataset.
In the realm of graph-based detection models, TextGCN, while performing well in text classification, showed mediocre results in code vulnerability detection experiments, achieving only a 91.50% accuracy rate on DiverseVul, with precision and recall at 15.66% and 11.50%, respectively. This may be due to TextGCN’s focus on word co-occurrence, lacking structural code information. The AST version of Devign, which incorporates control flow and data flow information and uses Word2Vec along with the average of tokens for node vector representation, performed poorly with an accuracy of only 70.21%, a gap of 14.26% - 23.31% from our F1 score. This could be attributed to the neglect of local semantic information of code within the node.
In vulnerability detection, false negatives (missed vulnerabilities) pose a significant risk. Vul-LMGNNs demonstrate strong recall across datasets, reducing false negatives compared to baseline models like Devign and ReVeal. On the ReVeal dataset, our Vul-LMGNN model achieved a recall of 51.41%, significantly higher than ReVeal’s 40.91%. In security applications, lowering false negatives is critical, and our model’s enhanced recall showcases its robustness in identifying vulnerabilities that other methods might miss. Additionally, we observed that false positives, although slightly higher in our model, are a reasonable trade-off for the improved vulnerability detection coverage.
The performance disparity on the other two datasets without specific CWEs is similar to those observed in the previous CWE-specific evaluations, as illustrated in Table 3. our proposed Vul-LMGNNs models across both the Devign and ReVeal datasets, showcasing significant improvements over existing approaches in vulnerability detection tasks. On the Devign dataset, Vul-LMGNNs(C-T5b) variant achieved the highest scores, with an accuracy of 67.27%, precision of 64.73%, recall of 62.20%, and F1 score of 63.44%. The performance gap is even more pronounced on the ReVeal dataset. Here, Vul-LMGNNs(C-T5b) attained remarkable results with an accuracy of 91.68%, precision of 54.89%, recall of 51.41%, and F1 score of 53.09%. The performance disparity between Vul-LMGNNs and other models is particularly evident when comparing to traditional approaches. For instance, on the ReVeal dataset, the best performing Vul-LMGNNs model achieved an F1 score that is nearly 17 percentage points higher than best baseline (53.09% vs. 36.48%).
In summary, the experimental results provide strong evidence for the effectiveness of Vul-LMGNNs in vulnerability detection tasks. The models’ ability to capture both structural and semantic information in code results in improved recall, lower false negatives, and consistent outperformance across datasets. The models’ ability to outperform existing state-of-the-art approaches across multiple metrics and datasets underscores their potential to significantly advance the field of automated code vulnerability detection.
5.2 Impact of the Tradeoff Parameter (RQ2)